# Heterogeneous Treatment Effects and

### v2

@@ 3 Framework
We introduce a framework that we use to analyze the relationship between HTEs—as estimated by treatment-covariate interactions—and the detection of causal mechanisms. Our framework is built upon the potential outcomes framework or Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974). We denote treatment by $Z$. In order to consider HTEs with respect to pre-treatment moderators, denote the vector of measured pre-treatment covariates by $X=(X_{1},X_{2},...,X_{K})\in\mathbb{R}^{K}$. Clearly, it need not be the case that all $K$ covariates moderate a given treatment effect.
@@ 3.1 Causal Effects
In order to understand when HTEs can allow for detection of mechanisms, we formalize existing informal conventions. To do so, we decompose the total effect (on a unit, $i$) into direct and indirect effects, as is standard in the causal mediation literature (Imai and Yamamoto, 2013). Let $M$ be the mediator of interest. We analyze the more general case of multiple mmediators in C. Given two treatment values, $z,z^{\prime}\in Z$, the total effect of $Z$ on $Y$ is:
@@
The direct effect, $DE(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediator $M$ at potential outcomes $M(z;X)$. Our preferred interpretation of the direct effect is the composite effect of all other mechanisms (aside from $M$). However, the direct effect could also include unmediated effects of treatment on an outcome. The indirect effect measures the effect on the outcome that operates by changing the potential outcome of mediator $M$. As is standard, we can then re-write the total effect as follows: [^9] Here, we assume that the direct and indirect effects do not vary at different levels of Z. See more discussion by Imai et al. (2011). All results in the paper hold with other decompositions. See additional discussions of general cases in the C and correlated mechanisms in the E.
@@ 3.2 Mechanisms and Outcomes
The characterization of directly- and indirectly-affected outcomes is distinct from whether an outcome is or could be observed by the analyst. As such, our definition is distinct from the concepts of latent and observed outcomes (e.g., Fariss, Kenwick, and Reuning, 2020). In principle, directly-affected and indirectly-affected variables could both be latent and/or observed. We illustrate these possibilities concretely in D using three examples from (hypothetical) studies with some type of learning or updating mechanism.
The distinction between directly- and indirectly-affected outcomes represents a theoretical commitment about the causal sequencing of outcomes. In contrast, the distinction between latent and observed outcomes represents both a theoretical commitment about the relationship between outcomes (though not necessarily the sequence) and empirical concerns about what outcomes are or could be observed by the researcher. As we will highlight, the distinction directly- and indirectly-affected matters for the use of HTEs to provide evidence of mechanism activation. When outcomes later in a sequence are produced by a (non-zero) linear function of previous outcomes, they can be treated as directly-affected outcomes. This accomodates popular linear rescalings like $Z$-score transformations of observed outcomes.
@@ 4 HTEs and Mechanisms
We then denote $\textbf{X}^{MDV}$ as the (possibly empty) set of all possible covariates that satisfy Definition 4 for the mechanism $M$ of a directly-affected outcome. Intuitively, if $X_{k}\in\textbf{X}^{MDV}$, then covariate $X_{k}$ can serve as an indicator for a mechanism/mediator of interest. [^12] A slightly stronger version of Definition 4 holds when $Y$ is continuously differentiable with respect to $M$ and $Z$ In this case, Definition 4 can be expressed as $\frac{\partial}{\partial X_{k}}\left(\frac{\partial Y}{\partial M}\frac{\partial M}{\partial Z}\right)\neq 0$. Under our definition of MDVs, it could be the case that $X_{k}$ moderates the effect of the treatment on the mediator. Interestingly, it could also be the case that $X_{k}$ moderates the effect of the mediator on the outcome. Both possibilities are depicted in Figure 2. The detection of mechanisms using treatment-by-covariate interactions requires researchers to postulate a $MDV$ for a given mechanism.
@@
However, postulating a $MDV$ is not sufficient to use HTEs to detect the activation of a mechanism. Specifically, we must also be concerned with whether a given covariate is a $MDV$ for other mechanisms (incorporated in the direct effect). Specifically, in order to do this, researchers need an exclusion restriction to limit the number of mediators that are affected by a given covariate, $X_{k}$. Logically, if a single covariate moderates a mechanism of interest and the direct effect, we cannot use HTEs in that covariate to isolate the mechanism of interest. Assumption 2 rules out the possibility that direct effects depend on $X_{k}$. In the more general multiple-mechanism case, we also need exclusion assumptions to rule out the possibility that the indirect effects produced by other mechanisms are moderated by $X_{k}$ (see C).
@@
Assumption 2 constrains the relationship between a moderator, $X_{k}$, and the direct effect (taken to be a composite of other possible mechanims). Figure 3 depicts a violation of Assumption 2 graphically for a MDV ($X_{k}$) of the mechanism $M$. Assumption 2 rules out the blue dashed path from $X_{k}$ to the direct effect of $Z$ on $Y$. Given our notation of interactions in a DAG, this means that $X_{k}$ does not moderate the direct effect of $Z$ on $Y$. This assumption is necessary for the use of HTEs to detect mechanisms. In its absence, we cannot link a heterogeneous treatment effects to a unique theoretical mechanism representation (here, $M$) because we have no way to ensure additive separability of the effect of $M$ from the effects of other mechanisms. Importantly, Assumption 2 does not imply that $X_{k}$ must be independent of the direct effect of $Z$ on $Y$; these quantities can be arbitraily correlated so long as $X_{k}$ does not induce heterogeneity in the direct effect.
@@
Use of HTEs in order to learn about mechanism activation represents one alternative to mediation analysis. Assumptions 2 forms the core assumption underpinning this use of HTE. The comparison to mediation analysis invites a comparison of the exclusion assumption to the assumption of sequential ignorability. It is useful to note that there is no logical ordering of the two types of assumptions: the exclusion assumption does not imply sequential ignorability, nor does sequential ignorability imply the exclusion assumption. This means that HTEs cannot said to be a “more agnostic” or “less agnostic” quantitative test of mechanism activation than mediation. [^13] Mediation analysis attempts to estimate the influence of different mechanisms by decomposing total effects into other causal estimands. When our primary interest is the activation of a mechanism (as in HTEs analysis), we can say that a mechanism is activated in mediation analysis if its indirect effect is non-zero. In some applications, one set of assumptions may be more plausible or defensible than the other, but we cannot make a general claim about the strength of these distinct sets of assumptions. We provide a broader discussion comparing the use of HTEs to mediation analysis in E.
@@ 4.1 HTEs and Mechanisms with Directly-Affected Outcomes
Proposition 1 conveys important implications about our ability to use heterogeneous treatment effects to provide evidence for a mechanism. Recall that if $X_{k}$ is a MDV for mechanism $M$, $AIE(X_{k}=x)\neq AIE(X_{k}=x^{\prime})$ for some $x,x^{\prime}\in X_{k}$. This provides evidence that mechanism $M$ is active for at least one unit. The exclusion assumption rules out the possibility that $ADE(X_{k}=x)\neq ADE(X_{k}=x^{\prime})$. When the assumption holds, HTEs in $X_{k}$ are thus sufficient to show that $X_{k}$ is a MDV for mechanism $M$. This conforms to standard interpretations that the presence of HTEs support arguments about mechanism activation. Nevertheless, this interpretation invokes the exclusion assumption, which is generally not invoked explicitly. We now consider the converse: the case when there exist no HTEs with respect to $X_{k}$.
@@
Comparing Propositions 1 and 2, under the exclusion assumption, the presence of HTEs provides more information with regard to mechanisms than does the absence of HTE. In this sense, relying upon the a lack of heterogeneity to “rule out” a potential mechanism requires much stronger theoretical assumptions than is generally acknowledged. Specifically, we would need to assume that $X_{k}\in\textbf{X}^{MDV}$ in order to rule out the activation of mechanism $M$. Indeed, such an assumption is precisely what we are trying to learn from the presence of HTEs in Proposition 1.
@@ 5 Indirectly-Affected Outcomes
When outcomes are directly affected by a mechanism under a theoretical model, the existence of HTEs provides evidence that a mechanism is active, as we have shown in Proposition 1. Yet, the situation is more complicated in the case of indirectly-affected outcomes. As in our discussion of the exogenous shocks and voting, for the vote choice outcome, we observed heterogeneous treatment effects in both prior beliefs (the mechanism in the model) and in valence (not the mechanism in the model). In this section, we show that this finding is general to situations in which indirectly-affected outcomes are generated by a non-linear transformation of the relevant directly-affected outcome, as noted in Definition 3.
@@
In this case, assumption 2 holds by construction. Now, suppose that a researcher mistakenly thought that $X_{2}$ was a MDV for a mechanism (either $M$ or a non-existent mechanism). Since we have constructed the data generating process, we know that it is not: $X_{2}\in\textbf{X}^{R}-\textbf{X}^{MDV}$. Evaluating the CATE of $Z$ on turnout when $X_{2}=1$, we have:
@@
While it is straightforward to construct infinte such examples numerically, it is reasonable to ask: is this plausible in “real world” examples? We cannot answer this question empirically without observing which mechanisms are at work. Indeed, this is the whole problem that quantitative analysis of mechanisms seeks to (indirectly) answer! However, we can conduct simulations that use real data. In H, we use 2020 ANES data in a Monte Carlo simulation motivated by the theoretical model of Little, Schnakenberg, and Turner (2022). This paper proposes two mechanisms that account for how citizens update beliefs in response to new information: accuracy and directional motives. Their model implies that partisanship (or ideology) should be a MDV for directional but not accuracy motives, which would allow researchers to look for heterogeneity in partisanship to assess the presence of directional motives as a mechanism.
Following this logic, we simulate different treatment effects on a measure of latent attitudes about greenhouse gas regulation. Some simulations allow for directional motives, others shut down this channel. We show that when treatment has a non-zero effect on the latent attitudes for any subset of partisans, as we would expect from accuracy motives alone, there are HTEs in partisanship on a binary measure of preferences for stronger greenhouse gas regulation. This means that we cannot use these HTEs to assess the presence of directional motives with this indirectly measured outcome. Further, using the ANES sample, for some simulated effect sizes, there is greater statistical power to detect heterogeneity in partisanship when simulated effects on the latent variable are homogeneou in partisanship than when they are heterogeneous.
@@
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumptions 1 holds. If HTEs do not exist with respect to $X_{k}$, then $_{k}\in\textbf{X}.$ X .
@@ 6.1 Three essential theoretical questions
Question #1 is fairly standard in applied empirical research. Researchers often posit one or more mechanisms of interest in addition to alternative explanations. In the exposition of our framework, we focus on one mechanism and relegate alternative explanations to the composite direct effect. C provides guidance for cases with multiple mechanisms of interest.
@@ 6.4 Strengthening assumptions to use HTE as mechanism tests for indirectly-affected outcomes
Our results suggest that HTEs or lack thereof are uninformative about mechanism activation when measured outcomes are indirectly affected by treatment. It is worthwhile to consider whether we can make progress in this common case by imposing stronger assumptions. To this end, we consider two alternatives. First, we consider an empirical assumption that is widely utilized in partial identification results: monotonicity (Manski, 1997). In our context, monotonicity holds that for all $x^{\prime}>x\in X_{k}$, $CATE(x^{\prime})\geq(\leq)CATE(x)$. Second, we consider the invocation of theoretical assumptions that provide a mapping between an unobserved directly-affected outcome and an observed indirectly-affected outcome.
First, in Appendix G.1, we show that the assumption of montonicity is not, in general, sufficient to provide information about mechanism activation through analysis of HTEs. This occurs because an assumption of montonicity fails to provide enough information about the data generating process that generates directly- and indirectly-affected outcomes. We show this in the context of the following oft-used data generating process for a binary outcome, $h(Y)\in\{0,1\}$:
@@
Second, and in contrast, Appendix G.2 shows how widely-used random utility models do permit inferences about mechanism activation when observed outcomes are indirectly-affected. These models provide a functional mapping between an individual’s utility and their choice between relevant alternatives. For example, in our motivating example, an individual voter’s utility from a given candidate is not generally observed whereas their self-reported vote choice is observed. A random utility model decomposes utility from a given choice into an observed systematic component and a random component. By specifying the systematic component as a function of individual- or choice-specific covariates and assuming that the random component is distributed according to a extreme value distribution, researchers can estimate the systematic component of utility. Using these estimates of our directly-affected outcome (utility), researchers can then use HTE to assess mechanism activation under Proposition 1. The invocation of a random-utility model is not free: it makes strong assumptions about the mapping between utility and choice outcomes that researchers may not be well-positioned to make or assess. However, these assumptions about the mapping between directly- and indirectly-affected outcomes do buy additional information which may permit learning about mechanism activation from HTE.

### v3

@@ 3 Framework
We introduce a framework that allows us to analyze the relationship between HTEs—as estimated by treatment-covariate interactions—and the detection of causal mechanisms. Our framework is built upon the potential outcomes framework or Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974). We denote treatment by $Z$. In order to consider HTEs with respect to pre-treatment moderators, denote the vector of measured pre-treatment covariates by $X=(X_{1},X_{2},...,X_{K})\in\mathbb{R}^{K}$. As in our running example, it need not be the case that all $K$ covariates moderate a given treatment effect: in that example, the prior ($\pi_{i}$) moderates the effect but valence ($v_{i}$) does not.
@@ 3.1 Causal Effects
In order to understand when HTEs permit detection of mechanisms, we formalize existing informal conventions. To do so, we decompose the total effect (on a unit, $i$) into direct and indirect effects, as is standard in the causal mediation literature (Imai and Yamamoto, 2013). Let $M$ be the mediator of interest. We analyze the more general case of multiple mmediators in Appendix C. Given two treatment values, $z,z^{\prime}\in Z$, the total effect of $Z$ on $Y$ is:
@@
The direct effect, $DE(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediator $M$ at potential outcomes $M(z;X)$. Our preferred interpretation of the direct effect is the composite effect of all other mechanisms (aside from $M$). However, the direct effect could also include unmediated effects of treatment on an outcome. In our running example, there is only one mechanism—voter learning—so the direct effect (all other mechanisms) is zero. This is evident because the treatment has zero effect if we fix the posterior belief to the same level. The indirect effect measures the effect on the outcome that operates by changing the potential outcome of mediator $M$. In our example, the indirect effect measures the effect that passes through the learning mechanism. With a single mechanism, the indirect effect is equivalent to the total effect of treatment.
As is standard, we can re-write the total effect as follows: [^10] Here, we assume that the direct and indirect effects do not vary at different levels of Z. See more discussion by Imai et al. (2011). All results in the paper hold with other decompositions. See additional discussions of general cases in the Appendix C and correlated mechanisms in the Appendix E.
@@ 3.2 Mechanisms and Outcomes
The characterization of directly- and indirectly-affected outcomes is distinct from whether an outcome is or could be observed by the analyst. Consequently, our definition is distinct from the concepts of latent and observed outcomes (e.g., Fariss, Kenwick, and Reuning, 2020). In principle, directly-affected and indirectly-affected variables could both be latent and/or observed. We illustrate these possibilities concretely in Appendix D using three examples from (hypothetical) studies with some type of learning or updating mechanism. In empirical studies akin to our runnning example, the indirectly-affected outcome, vote choice (or self-reported vote choice), is generally observed whereas the directly-affected outcome, voter utility, is not. But this reflects a choice by researchers. Rather than measuring only vote choice, a researcher could invest in eliciting willingness-to-pay for the incumbent in an endline survey. This investment would render a measure of the directly-affected outcome observed.
The distinction between directly- and indirectly-affected outcomes represents a theoretical commitment about the causal sequencing of outcomes. In contrast, the distinction between latent and observed outcomes represents both a theoretical commitment about the relationship between outcomes and empirical concerns about what outcomes are or could be observed by the researcher. As we will highlight, the distinction directly- and indirectly-affected matters for the use of HTEs to provide evidence of mechanism activation. When outcomes later in a sequence are produced by a (non-zero) linear function of previous outcomes, they can be treated as directly-affected outcomes. This accomodates popular linear rescalings like $Z$-score transformations of observed outcomes.
@@ 4 HTEs and Mechanisms
We then denote $\textbf{X}^{MDV}$ as the (possibly empty) set of all possible covariates that satisfy Definition 4 for the mechanism $M$ of a directly-affected outcome. Intuitively, if $X_{k}\in\textbf{X}^{MDV}$, then covariate $X_{k}$ can serve as an indicator for a mechanism/mediator of interest. [^13] A slightly stronger version of Definition 4 holds when $Y$ is continuously differentiable with respect to $M$ and $Z$ In this case, Definition 4 can be expressed as $\frac{\partial}{\partial X_{k}}\left(\frac{\partial Y}{\partial M}\frac{\partial M}{\partial Z}\right)\neq 0$. Under our definition of MDVs, it could be the case that $X_{k}$ moderates the effect of the treatment on the mediator. Interestingly, it could also be the case that $X_{k}$ moderates the effect of the mediator on the outcome. Both possibilities are depicted in Figure 2. In our running example, a variable measuring voters’ prior beliefs would be a MDV for the learning mechanism whereas a variable measuring valence assessments would not. This follows immediately from the DAG in Figure 1.
While theory—as in our running example—aids researchers in postulating MDVs for a given mechanism, empirical researchers seek to understand whether that mechanism is indeed active. Researchers then seek to detect evidence of a mechanism using treatment-by-covariate interactions with a proposed MDV. In other words, given a proposed MDV $X_{k}$, we want to learn whether $X_{k}\in\textbf{X}^{MDV}$.
@@
However, postulating a MDV is not sufficient to use HTEs to detect the activation of a mechanism. Specifically, we must also be concerned with whether a given covariate is a MDV for other mechanisms (and therefore incorporated in the direct effect). To this end, researchers need an exclusion restriction to limit the number of mechanisms that are moderated by a given covariate, $X_{k}$. Logically, if a single covariate moderates a mechanism of interest and the direct effect, a finding of HTEs in that covariate cannot isolate the mechanism of interest. Assumption 1 rules out the possibility that direct effects depend on $X_{k}$. In the more general multiple-mechanism case, we also need exclusion assumptions to rule out the possibility that the indirect effects produced by other mechanisms are moderated by $X_{k}$ (see Appendix C).
@@
Assumption 1 constrains the relationship between a moderator, $X_{k}$, and the direct effect (taken to be a composite of other possible mechanims). Figure 3 depicts a violation of Assumption 1 graphically for a MDV ($X_{k}$) of the mechanism $M$. Assumption 1 rules out the blue dashed path from $X_{k}$ to the direct effect of $Z$ on $Y$. Given our notation for interactions in a DAG, this means that $X_{k}$ cannot moderate the direct effect of $Z$ on $Y$. In the absence of Assumption 1, we cannot link a heterogeneous treatment effects to a unique theoretical mechanism representation (here, $M$) because we have no way to ensure additive separability of the effect of $M$ from the effects of other mechanisms. Importantly, Assumption 1 does not imply that $X_{k}$ must be independent of the direct effect of $Z$ on $Y$; these quantities can be arbitraily correlated so long as $X_{k}$ does not induce heterogeneity in the direct effect.
@@
Use of HTEs in order to learn about mechanism activation represents one alternative to mediation analysis. Assumptions 1 forms the core assumption underpinning this use of HTE. The comparison to mediation analysis invites a comparison of the exclusion assumption to the assumption of sequential ignorability. It is useful to note that there is no logical ordering of the two types of assumptions: the exclusion assumption does not imply sequential ignorability, nor does sequential ignorability imply the exclusion assumption. This means that HTEs cannot said to be a “more agnostic” or “less agnostic” quantitative test of mechanism activation than mediation. [^14] Mediation analysis attempts to estimate the influence of different mechanisms by decomposing total effects into other causal estimands. When our primary interest is the activation of a mechanism (as in HTEs analysis), we can say that a mechanism is activated in mediation analysis if its indirect effect is non-zero. In some applications, one set of assumptions may be more plausible or defensible than the other, but we cannot make a general claim about the strength of these distinct sets of assumptions. We provide a broader discussion comparing the use of HTEs to mediation analysis in Appendix E.
@@ 4.1 HTEs and Mechanisms with Directly-Affected Outcomes
Proposition 1 conveys important implications about our ability to use heterogeneous treatment effects to provide evidence for a mechanism. Recall that if $X_{k}$ is a MDV for mechanism $M$, $AIE(X_{k}=x)\neq AIE(X_{k}=x^{\prime})$ for some $x,x^{\prime}\in X_{k}$. This provides evidence that mechanism $M$ is active for at least one unit. The exclusion assumption rules out the possibility that $ADE(X_{k}=x)\neq ADE(X_{k}=x^{\prime})$. When the assumption holds, HTEs in $X_{k}$ are thus sufficient to show that $X_{k}$ is a MDV for mechanism $M$. This conforms to standard interpretations that the presence of HTEs support arguments about mechanism activation. Nevertheless, this interpretation invokes the exclusion assumption, which is generally not invoked explicitly. In our running example, we showed the existence HTE in the voters’ priors ($\pi_{i}$) for the directly-affected voter utility outcome. If one were to detect heterogeneity in an empirical study in this class, the heterogeneity would provide evidence that voter learning is an active mechanism.
We now consider the converse: the case when there exist no HTEs with respect to $X_{k}$.
@@
In our running example, recall that theoretically there do not exist HTEs in valence for the directly-affected voter utility outcome. Within the model, we know that the valence mechanism is inert. However, in an empirical exercise, the analogous finding of no detectable heterogeneity provides limited information. Proposition 2 and Corollary 1 enumerate three possibilities: we chose the wrong MDV for an active valence mechanism; the valence mechanism is inert (as in the theory); or there exists no MDV for an active valence mechanism. Crucially, the empirical finding of no (detectable) heterogeneity cannot distinguish between these possibilities. Therefore, we cannot “rule out” a mechanism due to a lack of observed HTEs.
Comparing Propositions 1 and 2, the presence of HTEs provides more information with regard to mechanisms than does the absence of HTEs. In this sense, relying upon the a lack of heterogeneity to “rule out” a potential mechanism would require even stronger assumptions. Specifically, we would need to assume that $X_{k}\in\textbf{X}^{MDV}$ in order to rule out the activation of mechanism $M$. Indeed, such an assumption is precisely what we are trying to learn from the presence of HTEs in Proposition 1.
@@ 5 Indirectly-Affected Outcomes
When outcomes are directly affected by a mechanism under a theoretical model, the existence of HTEs provides evidence that a mechanism is active, as we have shown in Proposition 1. Yet, the situation is more complicated in the case of indirectly-affected outcomes. As in our discussion of the exogenous shocks and voting, for the vote choice outcome, there exist HTEs in both prior beliefs (the mechanism in the model) and in valence (not the mechanism in the model). In this section, we show that this finding is general to situations with indirectly-affected outcomes as defined in Definition 3.
@@
The result and logic in Proposition 3 shows precisely why we see HTEs in both the priors and valence with the vote choice outcome in our running example. Even though valence is unrelated to the learning mechanism, it does enter the voter’s utility in (2). This means that a variable measuring valence would be relevant for the the directly-affected outcome. Yet, because vote choice is a non-linear function of this outcome, there exist—and we may observe—HTEs with respect to valence. This suggests that the qualitative existence HTEs with respect to an indirectly-affected outcome are not useful for the detection of mechanisms.
@@
In this case, Assumption 1 holds by construction since there is only one mechanism. Now, suppose that a researcher mistakenly thought that $X_{2}$ was a MDV for a mechanism (either $M$ or an inert mechanism). Since we have constructed the data generating process, we know that it is not: $X_{2}\in\textbf{X}^{R}-\textbf{X}^{MDV}$. Evaluating the CATE of $Z$ on turnout when $X_{2}=1$, we have:
@@
While it is straightforward to construct infinte such examples numerically, it is reasonable to ask: is this plausible in “real world” examples? Unfortunately, it is not possible to provide a purely empirical example because mechanisms are theoretical, not empirical objects. While we can, in principle, observe and measure the effects of a mechanism, we do not observe mechanisms themselves (Slough and Tyson, 2024a). [^16] The status of mechanisms as theoretical, not empirical, objects is precisely why mechanism testing is so difficult. However, simulations using real data and an assumed theory allow us to assess the likelihood that we observe heterogeneity when a covariate is or is not an MDV. In Appendix H, we use 2020 ANES data in a Monte Carlo simulation motivated by the theoretical model of Little, Schnakenberg, and Turner (2022). This paper proposes two mechanisms that account for how citizens update beliefs in response to new information: accuracy and directional motives. Their model implies that partisanship (or ideology) should be a MDV for directional but not accuracy motives, which would allow researchers to look for heterogeneity in partisanship to assess the presence of directional motives as a mechanism.
Following this logic, we simulate different treatment effects on a measure of latent attitudes about greenhouse gas regulation. Some simulations allow for directional motives, others shut down this channel. We show that when treatment has a non-zero effect on the latent attitudes for any subset of partisans, as we would expect from accuracy motives alone, there are HTEs in partisanship on a binary measure of preferences for stronger greenhouse gas regulation. This means that we cannot use these HTEs to assess the presence of directional motives with this indirectly measured outcome. Further, using the ANES sample, for some simulated effect sizes, there is greater statistical power to detect heterogeneity in partisanship when simulated effects on the latent variable are homogeneous in partisanship than when they are heterogeneous.
@@
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumption 1 holds. If HTEs do not exist with respect to $X_{k}$, then $_{k}\in\textbf{X}.$ X .
@@ 6.1 Three essential theoretical questions
Question #1 is fairly standard in applied empirical research. Researchers often posit one or more mechanisms of interest in addition to alternative explanations. In the exposition of our framework, we focus on one mechanism and relegate alternative explanations to the composite direct effect. Appendix C provides guidance for cases with multiple mechanisms of interest.
@@ 6.4 Strengthening assumptions to use HTE as mechanism tests for indirectly-affected outcomes
Our results suggest that HTEs or lack thereof are uninformative about mechanism activation when measured outcomes are indirectly affected by treatment. It is worthwhile to consider whether we can make progress in this common case by imposing stronger assumptions. To this end, we consider three alternatives. First, we consider an empirical assumption that is widely utilized in partial identification results: monotonicity (Manski, 1997). In our context, monotonicity holds that for all $x^{\prime}>x\in X_{k}$, $CATE(x^{\prime})\geq(\leq)CATE(x)$. Second, we consider the invocation of theoretical assumptions that provide a mapping between an unobserved directly-affected outcome and an observed indirectly-affected outcome. Finally, we ask whether moving from the existence to the magnitude of HTE can generate more information about mechanisms.
Monotonicity: First, in Appendix G.1, we show that the assumption of montonicity is not, in general, sufficient to provide information about mechanism activation through analysis of HTEs. This occurs because an assumption of montonicity fails to provide enough information about the data generating process that generates directly- and indirectly-affected outcomes. We show this in the context of the following oft-used data generating process for a binary outcome, $h(Y)\in\{0,1\}$:
@@
Modelling directly-affected outcomes: Second, and in contrast, Appendix G.2 shows how widely-used random utility models do permit inferences about mechanism activation when observed outcomes are indirectly-affected. These models provide a functional mapping between an individual’s utility (the directly-affected outcome) and their choice between relevant alternatives (the indirectly-affected outcome). For example, in our motivating example, a random utility model aims to recover a measure of utility by decomposing a given decision (i.e., vote choice) into an observed systematic component and a random component. By specifying the systematic component as a function of individual- or choice-specific covariates and assuming the distribution of the random component, researchers can estimate the systematic component of utility. Using these estimates of our directly-affected outcome (utility), researchers can then use HTE to assess mechanism activation under Proposition 1. The invocation of a random-utility model is not free: it makes strong assumptions about the mapping between utility and choice outcomes. Researchers may not be well-positioned to invoke or assess these assumptions. However, these assumptions about the mapping between directly- and indirectly-affected outcomes do buy additional information which may permit learning about mechanism activation from HTE.
Learning from the Magnitude of HTEs: To this point, following common practice, we have focused on the existence of HTEs, rather than the magnitude of HTEs. By magnitude of HTEs, we are referring to the the size of the difference in CATEs. In principle, this magnitude may provide additional information. For example, suppose that theory suggested that a given moderator $X_{k}$ may be a MDV for two distinct mechanisms. If this were the case, the exclusion assumptions (Assumptions 1 and 3) would not be satisfied. However, we may have reasons to believe that the HTEs coming from one mechanism are large (in magnitude) while those coming from the other mechanism are small (in magnitude). In Appendix G.3, we propose a simple Bayesian model that allows for inference about the activation of a specified mechanism from the magnitude of estimated effects. This model requires specification of priors over: (1) the probability of activation of a given mechanism and (2) the magnitude of HTEs under each candidate mechanism. Specification of these priors constitutes the invocation of additional assumptions.
We note that most theories in the social sciences admit directional—rather than point—predictions. By moving from the existence of HTE to their magnitude, the model that we propose requires researchers to specify priors about the size of HTEs, not simply their existence or direction. While these priors allow us to glean some information about mechanisms when the exclusion assumption(s) are violated, more work is needed to guide researchers in specifying such priors from theories that are largely directional.